Crystal is taking out an amortized loan for​ $16,000 to remodel parts of her home. Find the monthly payment if the loan is taken at​ 4% interest compounded monthly and she has agreed to pay it back over 8 years.
Answer by Theo(13342)

present value = 16,000
interest rate is 4% per year compounded monthly.
term of loan is 8 years.
formula to use is:
ANNUITY FOR A PRESENT AMOUNT WITH END OF TIME PERIOD PAYMENTS
a = (p*r)/(1-(1/(1+r)^n))
a is the annuity.
p is the present amount.
r is the interest rate per time period.
n is the number of time periods.
p = 16000
r = 4/1200
n = 8*12
formula becomes:
a = (16000*12/100)/(1-(1/(1+4/1200)^(8*12))) = 195.0284048
you can use a calculator as well.
i used the TI-BA-II
inputs were:
present value = 16000
interest rate per month = (4/12)%
number of months = 8*12
future value = 0
payments are made at the end of each time period.
click on payment to get payment at the end of each month = the same as with the formula.